# A Fully Adaptive Radau Method for the Efficient Solution of Stiff Ordinary Differential Equations at Low Tolerances

### v1

@@ Abstract
Radau IIA methods, specifically the adaptive order radau method in Fortran due to Hairer, are known to be state-of-the-art for the high-accuracy solution of highly stiff ordinary differential equations (ODEs). However, the traditional implementation was specialized to a specific range of tolerance, in particular only supporting 5th, 9th, and 13th order versions of the tableau and only derived in double precision floating point, thus limiting the ability to be truly general purpose for highly accurate scenarios. To alleviate these constraints, we implement an adaptive-time adaptive-order Radau method which can derive the coefficients for the Radau IIA embedded tableau to any order on the fly to any precision. Additionally, our Julia-based implementation includes many modernizations to improve performance, including improvements to the order adaptation scheme and improved linear algebra integrations. In a head-to-head benchmark against the classic Fortran implementation, we demonstrate our implementation is approximately 2x across a range of stiff ODEs. We benchmark our algorithm against several well-reputed numerical integrators for stiff ODEs and find state-of-the-art performance on several test problems, with a 1.5-times speed-up over common numerical integrators for stiff ODEs when low error tolerance is required. The newly implemented method is distributed in open source software for free usage on stiff ODEs.
@@ I Introduction
High-precision solving of ordinary differential equations (ODEs) is required in many disciplines, from controls and safety scenarios such as identifying potential asteroid collisions [1], to robustly resolving chemical reactions in astrochemistry which span more time scales than which 64-bit floating point numbers can account for [2]. There are many traditional scenarios where high precision is required as well, for example, in pharmacometrics it is a requirement during inference to solve the model to a much lower tolerance than what’s generally used in simulation (tolerances of around 1e-10) in order to improve the robustness of the fitting schemes [3]. This is due to the fact that calculating derivatives of ODE solutions incurs error from the forward pass and induced solver error from the extended equations for the forward/adjoint pass, making higher forward simulation accuracy a requirement in order to achieve accurate gradients and thus a stable optimization scheme [4] [5].
@@
I-A Runge-Kutta Methods
Runge-Kutta methods are generalized methods to numerically approximating solutions to first order differential equations (ODEs) [20]. They are collocation methods that use intermediate approximations of the solution to extrapolate the overall solution.
@@ I-A Runge-Kutta Methods
Such a problem can numerically be approximated at some future point in time $t$ using a Runge-Kutta method. The number of stages $s$ of a Runge-Kutta method denotes the number of nodes at which to take approximations of the solution. For an $s$ stage Runge-Kutta method with step size $dt$, the solution to $(1)$ at $t$ can be given as:
$y(t)=y_{0}+dt\cdot\Sigma_{i=1}^{s}{b_{i}k_{i}}$ (3)
@@
for $1\leq p\leq s$.
The constants $a_{i,j}$ for $1\leq i,j\leq s$, $b_{i}$ for $1\leq i\leq s$, and $c_{i}$ for $1\leq i\leq s$ are typically represented by a Butcher tableau, as shown below:
@@
$a$ is defined as the internal coefficients, $b$ is defined as the weights, and $c$ is defined as the nodes [7]. Throughout this paper, we take $c_{s}=1$ as is defined for a RadauIIA method.
We define the order of a Runge-Kutta method to be the maximum number $n$ such that the Taylor series expansion of $y(t_{0}+dt)$ and $y_{1}$ coincide for all terms from $1$ to $n$ [7].
The constants $a_{i,j}$, $b_{i}$, and $c_{i}$ typically satisfy some order conditions that make them more usable. For example, the first order conditions for a Runge-Kutta Method is
$\Sigma_{i=1}^{s}{b_{i}}=1$ (5)
While the second order condition is:
$\Sigma_{i=1}^{s}{b_{i}c_{i}}=\frac{1}{2}$ (6)
Such order conditions extend to third order conditions and further, and are important for testing whether a given method truly matches its assigned order. Higher order methods should approximate the solution faster as the timesteps shrink, so these order conditions are important in checking whether the method is of the appropriate order. The derivations of the above definitions can be found in [7].
Several further ”simplifying assumptions” are commonly used, for example:
$\Sigma_{j=1}^{i-1}{a_{ij}}=c_{i}$ (7)
but this condition is neither necessary nor sufficient for consistency. Many similar conditions exist; they merely serve to simplify the derivation of coefficients [7].
Another important property of Runge-Kutta methods to consider is their stability. The stability of a numerical method describes the tendency of local errors to either be amplified or diminished across time-steps. In other words, the stability properties of a Runge-Kutta method determines whether or not the method will converge globally as $dt$ goes to 0 [11].
The stability function of a Runge-Kutta method is found when the method is applied to an ODE of the form $y^{\prime}=\lambda y$. When Runge-Kutta methods use selected $dt$ that fall in the stability region, they prove to be highly accurate at approximating the true solution [11].
I-B Stiff ODEs
A subset of ODEs, known as stiff ODEs, pose particular challenges for most numerical integrators. A stiff ODE is one where there exists a dissonance between a slow and fast phase in the ODE, where the slow phase needs to be resolved while avoiding divergence in the fast phase.
As an example of a stiff ODE, consider the well known chemical reaction known as the Robertson equation [11]:
[…]
When translating this system into a system of linear ODEs, there will be a clear difference in the rates of equations 8 and 9, which will lead most Runge-Kutta methods struggle to handle the contrast between these two phases. An efficient numerical integrator will take as large time-steps as is possible to approach the solution quickly, but doing so in this case would cause the method to diverge from the true solution because of the rapid changes in equation 9. It is here where we see that stability of a numerical integrator is crucial, as having a wider range of options for the step-size allows the method to be more efficient on these systems [11]. As we will see in the next section, Radau methods posses stability properties necessary to overcome these challenges.
I-C Radau Methods
Radau methods are a class of implicit Runge-Kutta methods that are defined by quadrature formulas. An explicit Runge-Kutta method is one where the derivation of the current state is only done using previous states [11]. The simplest example is the explicit Euler method:
$y(t_{0}+h)=y_{0}+dt\cdot f(t_{0}+dt,y_{0})$ (11)
However, implicit Runge-Kutta methods use an implicit definition of the current state in addition to past states in deriving the current state, as seen below, in the implicit Euler method:
$y_{1}=y_{0}+dt\cdot f(t_{1},y_{1})$ (12)
When discussing the benefits of using Radau methods, it’s crucial to consider their stability, which is one of their prize-properties. For solving stiff ODEs, strong stability properties are crucial to ensure that the numerical method can handle the perturbations [11].
Radau methods have those stability properties. They are A-stable methods, meaning that when they are applied to an ODE of the form $\frac{dy}{dx}=qx$, where $q$ has a negative real part, the solutions will trend to 0 for all positive $dt$ [11]. Graphically, this property means that our method is stable on the entire left half of the complex plane [11]. That large stability region allows Radau methods to converge across a large range of time-steps, an important property while trying to solve stiff ODEs.
Furthermore a more crucial property for solving stiff ODEs is L-stability, which a method posses when it is both A-stable and has the stability function $g(dt)$ go to 0 as $dt\rightarrow\infty$ [11]. This property makes the method robust to the oscillations characteristic of stiff ODEs, and Radau methods also have this property. As a result, they are among the most efficient numerical integration methods for solving stiff ODEs.
Consider the below comparison between one of the most popular explicit methods, Dopri15, and RadauIIA5, a 5th order Radau method. The difference in performance is glaring. Radau, with its bigger stability region and more optimal properties, quickly converges on the solution. However, Dopri, struggling to resolve the perturbations in the equation, fluctuates around the true solution.
Figure 1: Comparison between different Runge-Kutta methods on a stiff ODE (source: [11])
In this paper, we describe our implementation of a Radau method that builds on the existing art, increasing the speed at which stiff ODEs can be solved.
II The State of Radau and our Developments
The most popular Radau software was developed by Ernst Hairer and published in his landmark paper “Stiff differential equations solved by Radau methods”. There, he laid out his implementation of a Radau method that autonomously switches between methods of 5th, 9th, and 13th order [19]. By changing the order of the method, his algorithm can more appropriately adapt to challenges in solving the stiff ODE. If a certain part of solving an ODE requires smaller time-steps, the method can switch to a more efficient higher-order method. Likewise, if a method can take larger time-steps in another region of the solve, a lower order method helps to streamline the computation. So, by having an adaptive order method, the method can adjust with the problem. This allows the method to optimize both solve time and accuracy.
However, there were some limitations to the existing algorithm:
@@ II The State of Radau and our Developments
In our new adaptive Radau method, we address both of the limitations of Hairer’s initial work, thus enhancing the performance of the method.
@@ III-A Programming Language Choice
III-B Tableau Elements
In building the new adaptive Radau method, we developed an algorithm that autonomously calculates the tableau for a method of any given order. The derivation comes from the order conditions and simplifying assumptions, as shown below [21]:
$c$ is defined as the roots of the Radau polynomial:
$\frac{\mathrm{d}^{s-1}}{\mathrm{d}x^{s-1}}x^{s-1}(x-1)^{s}$ (13)
${a_{i,j}}$ is then defined as:
$\Sigma_{j=1}^{s}{a_{i,j}c_{j}^{q-1}}=\frac{c_{i}^{q}}{q},i=1,2,\cdots,s,q=1,2,\cdots,s$ (14)
$b$ is simply defined as:
$b_{i}=a_{s,i}$ (15)
Mathematically, this derivation would be sufficient for our Radau method. However, in implementing the method in code, we transform our computations into a basis where $a^{-1}$ has a defined form, because $a^{-1}$ appears while solving a costly linear system. Knowing the form of $a^{-1}$ helps streamline computation.
III-C Transformation Matrix
To understand the transformation matrix $T$, it is first important to understand the form of $a^{-1}$. Because we only operate on methods with an odd number of stages, $a^{-1}$ always has one real eigenvalue and several complex-conjugate pairs of eigenvalues. For example, when $s=3$, below is $a^{-1}$ and its eigenvectors:
@@ III-C Transformation Matrix
Thus, applying our transformation matrix $T$ allows us to have a sparser structure of $a^{-1}$, which can be exploited to split the LU-factorizations and thus greatly decrease the computational cost of the linear algebra.
III-D Embedded Set Calculation
A second element that our script derives is an adaptivity set. This adaptivity set is used to vary the step-size of our method. In regions where there is a low error-estimate for our method, we can take larger time- steps so as to optimize our solve time, while if there is a higher error-estimate, we take smaller time-steps to prevent divergence of our method.
The calculation of our time-steps depends on an error-estimate that is calculated based on the difference between our method and a lower order ”embedded method”. This method shares the tableau coefficients $a$ and $c$ with the original method, while having a modified $b$ set that allows the embedded method to have a lower order. The calculation of the $b$ set is done using the order conditions.
By taking the difference between the value calculated by the embedded method and our Radau method, we have an estimate of the error of the Radau method. Then, our predictive controller uses that error estimate to determine the optimal size for the next time-step, as will be described in the next section.
Our embedded set is calculated by considering our method as a collocation method [27], which then gives us the following linear system of equations that is easily solved for $b$:
$\Sigma_{j=1}^{s}{b_{i}c_{i}^{k-1}}=\frac{1}{k},k=1,2,\cdots,s$ (16)
III-E Tableau Caching
In our implementation, we make use of a dictionary to store the tableaus. This helps to alleviate some computation time and avoids repeated re-derivation of commonly used tableaus. However, it remains true that our method can calculate the coefficients for a method of any order on-the-fly. Caching the common tableaus merely helps to speed up the process for those most frequently used orders, like 5th, 9th, and 13th. If the method chooses to go beyond those orders, it will derive the tableau autonomously, cache the tableau, and continue the numerical integration process.
III-F Step-Size Adaptivity
The algorithm makes use of the standard predictive controller and step-size selection algorithm.
The general idea is to use our error estimate to determine whether our current step-size is too big or small, and to scale the size accordingly. A high error estimate indicates that we should take smaller time-steps, and vice-versa for low error estimates.
The controller changes $dt$ based on the error. The number of iterations needed to converge is also factored into the algorithm, as this is indicative of the convergence rate. Mathematically, we calculate parameters $f$ and $expo$ to determine the scaling factor, $q$.
$f=\frac{(1+2*max\_iter)*\gamma}{num\_iters+2*max\_iter}$ (17)
$expo=\frac{1}{(adaptive\_order+1)}\$ (18)
$q=\frac{error^{expo}}{f}$ (19)
Where $\gamma$, $max\_iter$, and $num\_iters$ are parameters for a given iteration of our solve. Then, we find our scale factor $q$, and multiply $dt$ by this value to determine what the next time-step should be [28].
III-G Order Adaptivity
In addition to step size adaptivity, our method also builds on the order adaptivity in existing methods by implementing a novel order selection algorithm. Where the existing implementation uses the value of the divergence to determine whether to raise or lower the order [19], our algorithm chooses a different approach. However, the general scheme is similar.
The time-steps of a Radau method uses several Newton iterations while converging on the solution to that timestep. So, we use a metric based on the number of iterations that it takes the method to converge. If the method is taking several iterations to converge, this implies that a lower order method will be more efficient.
@@ III-G Order Adaptivity
Doing so allows us to combine step-size and order adaptivity. When we are taking small timesteps and need more iterations to converge, the algorithm chooses a higher order method, and vice versa. Thus, we optimize the solving of each section of the problem, thus optimizing the solve as a whole.
@@ IV-A Competing algorithms
BDF and Rosebrock methods cannot achive the same order of accuracy that Radau methods can. Therefore, while BDF and Rosenbrock methods tend to dominate solving stiff ODEs at high tolerances, Radau methods typically dominate the solution to low tolerances.
@@ IV-B Work-Precision Diagrams
We also consider benchmarks of our algorithm beyond the Float64 domain. In the following, we compare our AdaptiveRadau algorithm against one of the best stiff numerical integrators in Julia, Rodas5P, a newer high-order Rosenbrock method [17]. Both solutions are compared with a reference solution computed to tolerances of $10^{-24}$. The problem and solutions are computed with BigFloat precision.
@@ Acknowledgment
S.E. thanks the MIT PRIMES Computer Science program for connecting him with his mentor and giving him the chance to work on this research, and Dr. Chris Rackauckas and Oscar Smith for their continued support throughout the process.

### v2

@@ Abstract
Radau IIA methods, specifically the adaptive order Radau method in Fortran due to Hairer, are known to be state-of-the-art for the high-accuracy solution of highly stiff ordinary differential equations (ODEs). However, the traditional implementation was specialized to a specific range of tolerance, in particular only supporting 5th, 9th, and 13th order versions of the tableau and only derived in double precision floating point, thus limiting the ability to be truly general purpose for high fidelity scenarios. To alleviate these constraints, we implement an adaptive-time adaptive-order Radau method which can derive the coefficients for the Radau IIA embedded tableau to any order on the fly to any precision. Additionally, our Julia-based implementation includes many modernizations to improve performance, including improvements to the order adaptation scheme and improved linear algebra integrations. In a head-to-head benchmark against the classic Fortran implementation, we demonstrate our implementation is approximately 2x across a range of stiff ODEs. We benchmark our algorithm against several well-reputed numerical integrators for stiff ODEs and find state-of-the-art performance on several test problems, with a 1.5-times speed-up over common numerical integrators for stiff ODEs when low error tolerance is required. The newly implemented method is distributed in open source software for free usage on stiff ODEs.
@@ I Introduction
High-precision solving of ordinary differential equations (ODEs) is required in many disciplines, from controls and safety scenarios such as identifying potential asteroid collisions [1], to robustly resolving chemical reactions in astrochemistry which span more time scales than which 64-bit floating point numbers can account for [2]. There are many traditional scenarios where high precision is required as well, for example, in pharmacometrics it is a requirement during inference to solve the model to a much lower tolerance than what’s generally used in simulation (tolerances of around $10^{-10}$) in order to improve the robustness of the fitting schemes [3]. This is due to the fact that calculating derivatives of ODE solutions incurs error from the forward pass and induced solver error from the extended equations for the forward/adjoint pass, making higher forward simulation accuracy a requirement in order to achieve accurate gradients and thus a stable optimization scheme [4] [5].
@@
II Background
II-A Fully-Implicit Runge-Kutta Methods
@@
For an $s$ stage Runge-Kutta method with step size $dt$, the approximate solution to $(1)$ at $t$ can be given as:
$y_{n+1}=y_{n}+dt\cdot\Sigma_{i=1}^{s}{b_{i}k_{i}}$ (3)
@@
for $1\leq p\leq s$, where the constants $a_{i,j}$ for $1\leq i,j\leq s$, $b_{i}$ for $1\leq i\leq s$, and $c_{i}$ for $1\leq i\leq s$ are typically represented by a Butcher tableau [7], as shown below:
@@
Explicit Runge-Kutta methods are methods for which the coefficient matrix $a_{i,j}$ is lower triangular. However, it is a well-known result that such methods have strict stability limits and thus implicit tableaus are required for systems with large condition numbers on the Jacobian, also known as stiffness [11].
II-B Radau IIA Methods
Radau methods are a class of implicit Runge-Kutta methods that are defined by the Radau quadrature formulas. The tableau is defined by the elements of the Radau polynomial [21], where $c$ is defined as the roots of the Radau polynomial:
$\frac{\mathrm{d}^{s-1}}{\mathrm{d}x^{s-1}}x^{s-1}(x-1)^{s}$ (5)
We then utilize the so-called ”simplifying assumptions” [21] to define two $s$ by $s$ matrices $c_{powers}$ and $c_{q}$ based on $c$, with each element defined as:
$c_{powers}[i][j]=c[i]^{j-1}$ (6)
$c_{q}[i][j]=c_{powers}[i][j]\cdot c[i]/j$ (7)
Then, the constants $a$ are defined as:
$a=c_{q}\cdot c_{powers}^{-1}$ (8)
$b$ is simply defined as:
$b_{i}=a_{s,i}$ (9)
as required by the order conditions. It can be shown that the $s$ stage Radau method gives an order of $12s-$ 1 and these methods are A-stable, L-stable, and B-stable for all orders [19], meaning that all Radau IIA methods generally have higher order than the number of stages and achieve the highest level of stability that is generally provable. It is this high order combined with the with stability that suggests their theoretical potential for highly-accurate integration of stiff systems.
II-C Hairer’s Fortran Radau Implementation
The most popular Radau software was developed by Ernst Hairer and published in his landmark paper “Stiff differential equations solved by Radau methods” [7]. There, he laid out his implementation of a Radau method that autonomously switches between methods of 5th, 9th, and 13th order [19]. An accompanying implementation in Fortran, simply known as radau, is a standard integrator used in all sorts of platforms from SciPy [22] to OpenModelica [23]. However, there were some limitations to the existing algorithm:
@@ III-A Programming Language Choice
III-B Arbitrary Order Tableau Generation and Caching
A notable departure from the previous Radau IIA implementations is the software directly uses the equation 8 for the derivation of the tableaus. The derivation of $a_{i,j}$ is generated using the precision of the provided $u(0)$ initial condition vector and the $c_{i}$ is generated using the precision of the chosen $(t_{0},t_{f})$ time span description. The computed tableaus are stored in a dictionary to cache the common tableaus and speed up the setup process for those most frequently used orders, like 5th, 9th, and 13th. If the method chooses to go beyond those orders, it will derive the tableau autonomously, cache the tableau, and continue the numerical integration process.
III-C Embedded Method Generation and Step-Size Adaptivity
In order to make a Runge-Kutta method adaptive, an alternative reducer $\tilde{b}_{i}$ vector is typically generated where:
$\tilde{y}_{n+1}=y_{n}+dt\cdot\Sigma_{i=1}^{s}{\tilde{b}_{i}k_{i}}$ (10)
is a lower order approximation. This lower order method is then used to derive an error estimate
$E=\|\tilde{y}_{n}-y_{n}\|$ (11)
which is then used as an approximation for the location truncation error for the adaptive step size controller with a standard predictive controller [7]. Because the embedded Runge-Kutta method to a Radau IIA method is not uniquely defined, we have to make a choice as to the lower order approximation. Following the choice made by Hairer and justified by others [28], we choose the following linear system of equations that is easily solved for $\tilde{b}$ and gives us a Runge-Kutta method of order $s$:
$\Sigma_{j=1}^{s}{\tilde{b}_{i}c_{i}^{k-1}}=\frac{1}{k},k=1,2,\cdots,s$ (12)
III-D Optimizing LU Factorizations via $a$ Transformation
In order to achieve a major speedup over naive implementations, we use a eigen-decomposition trick first exploited in the Hairer radau [7]. To understand the transformation matrix $T$, it is first important to understand the form of $a^{-1}$. Because we only operate on methods with an odd number of stages, $a^{-1}$ always has one real eigenvalue and several complex-conjugate pairs of eigenvalues. For example, when $s=3$, below is $a^{-1}$ and its eigenvectors:
@@ III-D Optimizing LU Factorizations via $a$ Transformation
Thus, applying our transformation matrix $T$ allows us to have a block-diagonal structure of $a^{-1}$. Thus when moving the Newton solving to the transformed space $z=Tu$, the LU factorizations within the Newton method decouple into $s-1$ $2n\times 2n$ blocks and one $n\times n$ block. Notably, our implementation allows for parallelizing these LU factorizations which can be an optimization when the blocks are too small for BLAS-based multithreading to be effective. However, this transformation is a major performance optimization even without parallelization as the decoupling of the $\mathcal{O}((sn)^{3})$ cost to LU factorize the $sn\times sn$ Newton matrix is transformed into $\mathcal{O}(sn^{3})$ as each block is done independently. This optimization is thus crucial to our implementation which can achieve high order and thus large stage sizes, effectively decreasing the cost by orders of magnitude for higher fidelity integrations.
III-E Order Adaptivity
In addition to step size adaptivity, our method also builds on the order adaptivity in existing methods by implementing a novel order selection algorithm. Where the Hairer implementation uses the value of the divergence to determine whether to raise or lower the order [19], our algorithm chooses a different approach. However, the general scheme is similar. The goal of the order adaptivity is to choose the highest order to get the largest time steps for which the Newton method can converge, and then the time step is chosen to be as large as possible in order to keep the error below tolerance. The time-steps of a Radau method uses several Newton iterations while converging on the solution to that timestep, where more iterations implies that the Newton solver is becoming unstable (and thus smaller time steps are required to get a better initial guess) and each iteration requires a fixed cost. Notably, this cost is only linear with respect to the order because of the eigenvector transformation of Section III-D, and thus the cost is both linear with respect to the chosen order and with respect to the number of Newton iterations. Thus we use a metric based on the number of iterations that it takes the method to converge. If the method is taking several iterations to converge, this implies that the Newton solver is at the edge of stability and that a lower order method can be more efficient.
@@ III-E Order Adaptivity
Note that increments of 4 are used so that every chosen order has the eigenvector structure required in Section III-D. Doing this approach allows us to combine step-size and order adaptivity. When we are taking small timesteps and need less iterations to converge, the algorithm chooses a higher order method, and vice versa.
@@ IV-A Competing algorithms
    SUNDIALS BDF Methods: A common C++ implementation of BDF methods. Chosen as a common “standard candle” of stiff ODE solvers, as CVODE is a standard method which is also very similar to other methods such as VODE and LSODE.
  4. 4.
@@
BDF and Rosebrock methods cannot achieve the same order of accuracy that Radau methods can. Therefore, while BDF and Rosenbrock methods tend to dominate solving stiff ODEs at high tolerances, Radau methods typically dominate the solution to low tolerances.
@@ IV-B Work-Precision Diagrams
We also consider benchmarks of our algorithm beyond the Float64 domain. In the following, we compare our AdaptiveRadau algorithm against one of the best stiff numerical integrators in Julia, Rodas5P, a newer high-order Rosenbrock method [17]. Both solutions are compared with a reference solution computed to tolerances of $10^{-24}$. The problem and solutions are computed with BigFloat precision. Notably, the difference in performance dramatically increases as the required tolerance increases, showcases the special high performance of this new integrator in this regime.
@@ Acknowledgment
S.E. thanks the MIT PRIMES Computer Science program for connecting him with his mentor and giving him the chance to work on this research, and Dr. Chris Rackauckas and Mr. Oscar Smith for their continued support throughout the process.